OpenStudy (elise_a18):

The triangle below has a total perimeter of 450 cm. : (triangle with sides 128 and 150. Third side unknown) Write an equation and solve it to find the length of the longest side of the triangle. 128x + 150 = 450; x = 2.34 cm x + 278 = 450; x = 172 cm 2x + 278 = 450; x = 86 cm 150x + 128 = 450; x = 2.15 cm

OpenStudy (anonymous):

128 + 150 + x = 450.

OpenStudy (anonymous):

278+x=450 x=450-278 x=?

OpenStudy (anonymous):

solve it @Elise_a18

OpenStudy (anonymous):

Then compare this answer with the given answer in the question

OpenStudy (triciaal):